OpenStudy (anonymous):

You deposit $20 at the end of every three months into an annuity with 4% interest compounded quarterly. How much of the annuity's value is from interest after 25 years?

OpenStudy (anonymous):

Use the Future Value of an Ordinary Annuity equation: FVoa = PMT [((1 + i)^n - 1) / i], where i = the quarterly rate 0.04/4 = 0.01, n = 25 yrs * 4 qtrs per year = 100, PMT is the quarterly deposit of $20... FVoa = 20 [(1.01^100) - 1 / 0.01] = 20 [(2.70481 - 1) / 0.01] = 20 [170.48138] = 3,409.63....now subtract your payments to get the amount of interest:... (20 * 100) = 2,000..interest is 1409.63, round to 1410

OpenStudy (anonymous):

Thank you so much ! (;